# Investigating Imperceptibility of Adversarial Attacks on Tabular Data: An Empirical Analysis

### v1

@@ 4.3 Evaluation of Attack Imperceptibility using Quantitative Properties
As mentioned before, an adversarial attack should achieve a success rate above a give threshold before its imperceptibility can be evaluated. We set a threshold of 30% for attack success rate. As shown in Table 2, applying the C&W attack to the LinearSVC model resulted in a success rate below the specified threshold across all five datasets. Therefore, it is excluded from the imperceptibility evaluation. Tables 4 to 8 present the results of imperceptibility of evaluation using four quantitative properties.
@@
Table 4: Results of sparsity of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models (LR, MLP and LinearSVC) over five datasets (Adult, German, COMPAS, Diabetes, and Breast Cancer). Best sparsity for each model is shown in bold. Due to unsatisfied effectiveness, the combination of C&W attack and LinearSVC is not considered and marked with $\dagger$.
@@
Sparsity metric quantifies the average number of modified features over all adversarial examples generated by an attack. Lower sparsity values suggest that the adversarial attacks possess better sparsity, implying that the modification made to the features are less perceptible. Since the number of features varies across datasets, sparsity can only be compared within the same dataset. The results in Table 4 suggest that all attack methods yield comparable sparsity results cross different models in each dataset.
Furthermore, we investigate the potential relationship between sparsity value and feature types. In two numerical feature only datasets, we observe that the sparsity value is consistently close to the total number of numerical features in each corresponding dataset (refer to Figure 1). The results imply that five attack methods exhibit a tendency to perturb almost all features within two numerical datasets. When analysing mixed datasets, comparing the sparsity value solely with the number of numerical or categorical features is insufficient to identify the relationship between them. We then examine all original inputs and their corresponding adversarial examples in each combination of attack and models. We calculate the individual feature perturbation count for each combination, which record the total number of times being perturbed by each feature. The results are visualised as heatmaps in Figure 2, which indicate that all attacks tend to alter numerical features more frequently than categorical features in Adult, German and Compas datasets. Only in Adult and German datasets, DeepFool and C&W attacks modify categorical features.
@@
Figure 2: The heatmaps of individual feature perturbation count for each individual feature in mixed datasets across different combination of models and attacks. Overall, four attacks almost perturb numerical feature only in three mixed datasets. Adversarial examples generated by DeepFool and C&W in Adult and German dataset are occasionally modified with categorical features. Due to unsatisfied effectiveness, the combination of C&W attack and LinearSVC is not considered.
@@
Proximity measures the similarity of the perturbed example and original input. A smaller proximity metric value indicates that adversarial examples are closer to original inputs, which represents better imperceptibility. In our evaluation, we use $\ell_{2}$ and $\ell_{\infty}$ norm as our two proximity metrics. The results of two metrics are shown in Table 5 ($\ell_{2}$) and Table 6 ($\ell_{\infty}$) respectively.
Table 5 suggests, in most datasets, C&W attack method yields the lowest Proximity $\ell_{2}$ values on both LR and MLP models, with the exception of German dataset and LR model. On the other hand, DeepFool achieves the lowest $\ell_{2}$ values using the LinearSVC model in all five datasets.
Table 6 suggests the proximity $\ell_{\infty}$ values for FGSM and PGD attacks are consistently 0.3 across all models and datasets. This is expected since bounded attacks are constrained to generate perturbations within this budget. As for the other three unbounded attacks, the $\ell_{\infty}$ norm values are influenced by different datasets, but no consistent patterns are observed. For instance, in Diabetes dataset, $\ell_{\infty}$ values of DeepFool attack are less than 0.3, but for other datasets, they may exceed it. In addition, three implemented models also have a potential impact on $\ell_{\infty}$ values, though results are inconsistent. For example, while DeepFool consistently achieves lower $\ell_{\infty}$ values on LinearSVC compared to LR and MLP, LowProFool does not exhibit the same trend.
Table 5: Results of proximity $\ell_{2}$ of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models (LR, MLP and LinearSVC) over five datasets (Adult, German, COMPAS, Diabetes, and Breast Cancer). Best $\ell_{2}$ value for each model is shown in bold. Due to unsatisfied effectiveness, the combination of C&W attack and LinearSVC is not considered and marked with $\dagger$.
@@
Table 6: Results of proximity $\ell_{\infty}$ of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models (LR, MLP and LinearSVC) over five datasets (Adult, German, COMPAS, Diabetes, and Breast Cancer). Best $\ell_{\infty}$ value for each model is shown in bold. Due to unsatisfied effectiveness, the combination of C&W attack and LinearSVC is not considered and marked with $\dagger$.
@@
The deviation metric assesses how perturbed examples relate to the overall dataset distribution. When the deviation metric is notably high, adversarial attacks are more likely to produce examples outside of the dataset’s distribution. Therefore, lower deviation values are preferable for crafting imperceptible attacks. In Table 7, the examination of the deviation metric reveals that C&W attack consistently results in the lowest deviation value, implying a higher degree of similarity to the original statistical distributions, especially in LR and MLP models. DeepFool, on the other hand, excels when applied to the LinearSVC model, achieving the best results in terms of deviation. The deviation level of LowProFool is lower than that of C&W and DeepFool attacks, but higher than two bounded attacks. In contrast, two bounded attacks, FGSM and PGD, consistently yield higher deviation values, suggesting that they tend to produce more outliers.
Table 7: Results of deviation of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models (LR, MLP and LinearSVC) over five datasets (Adult, German, COMPAS, Diabetes, and Breast Cancer). Best deviation for each model is shown in bold. Due to unsatisfied effectiveness, the combination of C&W attack and LinearSVC is not considered and marked with $\dagger$.
@@
Sensitivity quantifies how much narrow distribution features are altered by attack algorithms. Lower sensitivity values are preferable. No distinguishable patterns in sensitivity results have been observed crossing different attacks (see Table 8). In most combinations of models and datasets, C&W attack demonstrates relatively low values in sensitivity metric. DeepFool, LowProFool, FGSM, and PGD attacks consistently result in significantly higher sensitivity values. This indicates that these four attacks tend to introduce more significant perturbations on narrow distribution features than C&W.
Table 8: Results of sensitivity of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models (LR, MLP and LinearSVC) over five datasets (Adult, German, COMPAS, Diabetes, and Breast Cancer). Best sensitivity for each model is shown in bold. Due to low performance, the combination of C&W attack and LinearSVC is not considered and marked with $\dagger$.
@@ 4.4 Analysis of Imperceptibility using Qualitative Properties
In Section 3, we define three qualitative properties of attack imperceptibility, which are Immutability, feasibility and feature interdependency, and they often rely on domain-specific knowledge about the datasets. We analyse these qualitative properties using case-based examples.
@@
We analyse immutability across three mixed datasets (Adult, COMPAS, and German). Using common knowledge, we can identify examples of immutable features in each of these datasets as shown below:
@@
    German: Personal Status/Sex and Foreign Worker Status
We examine each adversarial example from these datasets to check whether the corresponding feature values have been modified. We can therefore assess whether the identified immutable features are susceptible to manipulation. We observe that DeepFool attempts to perturb immutable features in both LR and MLP models for all three dataset, clearly deviating from the immutability requirement. For instance, in the Adult dataset, DeepFool generates a total of 6,513 adversarial examples with the MLP model. In these examples, the feature Race is modified in 648 instances, Sex in 345 instances, and Marital Status in 1,096 instances. The results of perturbing immutable features for three datasets are detailed in A.1. Conversely, C&W, FGSM and PGD do not alter these immutable features. However, this does not inherently imply that three attacks adhere to the immutability requirement. Theoretically, these attacks do not take into account the perturbation of immutable features in their algorithm designs. Hence, they may not have capability to change these immutable features. This limitation arises because all identified immutable features are categorical variables that have been preprocessed using one-hot encoding for implementation. For two bounded attacks, their attack budgets are limited to 0.3, meaning they lack sufficient perturbation allowances to flip encoded categorical features. While C&W attacks is in unbounded attack setting, but it is not suitable to perturb encoded categorical features.
Here, we use the combination of COMPAS dataset and LR model as example to explain how Deepfool perturbs immutable features. COMPAS dataset is well-known tabular dataset with racial bias for scoring criminal defendant’s likelihood of recidivism (Barenstein, 2019), in which both feature race and feature sex are the immutable features. In our experiment, we classify each case as ”Medium-Low” or ”High” risk. We generate 1408 adversarial examples for DeepFool attack with LR model, and filtered cases with perturbation on immutable features. Two example cases (#285 and #501) shown in Table 9 are originally predicted as “Medium-Low”. DeepFool attack successfully misclassified these cases’ predictions, but notably, it also changed the individuals’ ethnicity from Caucasian and Hispanic to Native American and African-American.
Table 9: Comparison for differences of partial feature values between original datapoints and adversarial examples generated by DeepFool attack using Case #285 and Case #501.
@@
#285 No Attack 80 0 0 Greater than 45 Male Caucasian Medium-Low
@@
#501 No Attack 83 0 0 Greater than 45 Male Hispanic Medium-Low
@@
To understand the perturbations on immutable features, we analysed the feature weights (coefficients) of the LR model, as depicted in Figure 3(a). Notably, different values of feature race lead to different prediction contributions. As a result, DeepFool successfully perturbed the immutable feature race from categories with either positive or low negative impact (e.g., Caucasian and Hispanic) to categories with high negative impact (e.g., Native American and African-American).
(a) The feature weights of LR for COMPAS dataset after one-hot encoding: positive features contribute to “Medium-Low” risk and negative features contribute to “High” risk. Features with weights in the range of $[-0.15,0.15]$ have been omitted.
(b) The feature weights of LR for diabetes dataset: positive features contribute to “has” diabetes and negative features contribute to “not have” diabetes.
Figure 3: Feature weights of LR for COMPAS and diabetes datasets.
@@
Property feasibility assesses if adversarial attack algorithms can perturb feature values within a reasonable range. However, since there’s a shortage of background knowledge for all datasets, it’s impractical to identify all feasible feature ranges for each dataset. To assess the feasibility of adversarial attacks, we turn to the diabetes dataset, which exclusively comprises numerical features used to predict diabetes based on diagnostic measurements. Even though the original dataset includes descriptions of each variable (Smith et al., 1988), it does not specify feasible feature ranges. However, drawing on common medical knowledge, we have identified suitable ranges for certain features and have listed them in A.2.
Table 10: Comparison for differences of partial feature values between original datapoints and adversarial examples generated by DeepFool, C&W attack and LowProFool attack using Case #19 and Case #57.
@@
From this dataset, we’ve chosen two cases, labeled as #19 and #57, from the generated adversarial examples with LR models. We compare the original feature values with the adversarial examples generated by DeepFool, C&W attack, and LowProFool attack in Table 10 as following:
@@
    Glucose: For both Case #19 and Case #57, all three attack methods significantly elevate the values of the features Glucose from 100 to more than 135. The C&W attack introduces comparatively minor perturbations, maintaining Glucose values within the normal range (below 140). Conversely, the DeepFool and LowProFool attacks induce more substantial changes, pushing the glucose levels into the impaired glucose tolerance range (140–199). Particularly with LowProFool, it alters the Glucose values to reach 199 in both cases, which is extremely close to the diabetes threshold (200).
@@
    SkinThickness: The reference value range for women’s triceps skin fold thickness is 23.6$\pm$7.5mm. Although the adversarial examples generated by DeepFool (Case #19) and LowProFool (Case #19 & #57) exceed the maximum reference value, the deviation is not substantial.
@@
    Insulin: In Case #57, both the DeepFool and LowProFool attacks decrease the insulin level to below the lower bound of the recommended 2-hour serum insulin range (16–166). This perturbation is not consistent with expected physiological behavior.
@@
    BMI: Similar with In both cases, the BMI values are modified from around 30 to extremely obese levels, surpassing 40. Particularly in Case #57, the DeepFool and LowProFool attacks perturb the BMI to 59.85 and 67.10, respectively, which are unrealistic and uncommon scenarios.
The feature weights of LR model for diabetes dataset is extracted and visualised in Figure 3(b). It underscores that features Glucose and BMI exert a significantly positive impact on predicting diabetes patients. This explains why the three attacks prioritise these two features during perturbation, leading to changes in their values that extend into infeasible ranges.

### v3

@@ 4.3 Evaluation of Attack Imperceptibility using Quantitative Properties
As mentioned before, an adversarial attack should achieve a success rate above a reasonable threshold before its imperceptibility can be evaluated. We set a threshold of 30% for attack success rate. As shown in Table 4, applying the C&W attack to the LinearSVC model results in a success rate below the specified threshold across all five datasets. Therefore, the combination of the C&W attack on LinearSVC model is excluded from the imperceptibility evaluation. Tables 5 to 9 present the results of imperceptibility evaluation using four quantitative properties.
@@
Sparsity metric quantifies the average number of modified features over all adversarial examples generated by an attack. A lower sparsity value indicates that an adversarial attack modifies fewer features, suggesting an intention to make the perturbed inputs less noticeable. Hence, lower sparsity suggests better imperceptibility of an adversarial example. For tabular datasets, because the number of features is specific to each dataset, it is meaningful to compare sparsity only within the same dataset, not across different datasets. As shown in Table 5, the sparsity results indicate that all attack methods yield comparable sparsity cross different models within each dataset. In addition, the two bounded attacks, FGSM and PGD, often yield the same or very similar sparsity values.
Table 5: Results of sparsity of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models over five datasets. The best (i.e., lowest) sparsity for each model is shown in bold. Note that due to insufficient effectiveness, the combination of C&W attack on LinearSVC model is not considered and their results are marked with $\dagger$.
@@
We also investigate the potential relationship between sparsity and feature types. In the two numerical-only datasets (Diabetes and Breast Cancer), we observe that the sparsity value is consistently close to the respective total number of (numerical) features—8 for Diabetes and 30 for Breast Cancer (refer to Table 1). The results suggest that the five attack methods exhibit a tendency to perturb almost all features within these two datasets. When analysing the three mixed datasets (Adult, German and COMPAS), we first identify which features have been perturbed by comparing the original model inputs and their corresponding adversarial examples for each attack-model combination, and then count the total number of times each feature has been perturbed. Due to the varying test sizes of the evaluated datasets, these attacks produce different numbers of adversarial examples across these datasets. The results are visualised as heatmaps in Figure 2, suggesting that the four attacks (DeepFool, C&W, FGSM and PGD) primaly alter numerical features with minimal or no changes to categorical features in the three mixed datasets.
@@
Figure 2: Three heatmaps visualising the total count of each individual feature being perturbed across different attack/model combinations over three mixed datasets, respectively. The X axis enumerates all the features—numerical features followed by categorical features in each dataset—while the Y axis enumerates all the attack/model combinations. The colour scale of each plot is determined by the number of generated adversarial examples in each individual dataset: 6,513 examples for the Adult dataset, 1,443 examples for the COMPAS dataset, and 200 examples for the German dataset. Overall, the four attacks (DeepFool, C&W, FGSM and PGD) primarily perturb numerical features with minimal or no changes to categorical features in the three mixed datasets. Due to insufficient effectiveness, the combination of the C&W attack and LinearSVC is not considered.
@@
Proximity measures the similarity between the perturbed example and its original input. A smaller proximity metric value indicates that the adversarial example are closer to the original input, suggesting greater imperceptibility. In our evaluation, we use $\ell_{2}$ and $\ell_{\infty}$ norm as our two proximity metrics. We refer to them as proximity $\ell_{2}$ and proximity $\ell_{\infty}$, and the results of the two metrics are shown in Table 6 (proximity $\ell_{2}$) and Table 7 (proximity $\ell_{\infty}$), respectively.
From Table 6, the following observation can be made. Overall, the three unbounded attacks, C&W, DeepFool and LowProFool, yield smaller proximity $\ell_{2}$ values. C&W results in the lowest $\ell_{2}$ values on both LR and MLP models, with the exception of attacking the LR model on German dataset. When attacking the LinearSVC model, DeepFool achieves the lowest $\ell_{2}$ values across all datasets. The results of LowProFool exhibit higher $\ell_{2}$ values than C&W and DeepFool. The two bounded attacks, FGSM and PGD, have similar $\ell_{2}$ values on all models within the same dataset. The smaller $\ell_{2}$ values of the unbounded attacks suggest their ability to find minimal perturbation without attack budget constraints, while the similarity in $\ell_{2}$ values of the bounded attacks (within the same dataset) indicates that the uniform impact of their attack budget constraints is more significant than the influence of different predictive models.
Table 6: Results of proximity $\ell_{2}$ of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models over five datasets The best proximity $\ell_{2}$ value for each model is shown in bold. Due to nsufficient effectiveness, the combination of C&W attack and LinearSVC is not considered and their results are marked with $\dagger$.
@@
Table 7: Results of proximity $\ell_{\infty}$ of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive modelsover five datasets. The best proximity $\ell_{\infty}$ value for each model is shown in bold. Due to insufficient effectiveness, the combination of C&W attack and LinearSVC is not considered and their results are marked with $\dagger$.
@@
Table 7 shows that the proximity $\ell_{\infty}$ values for both FGSM and PGD attacks are consistently 0.3 across all models and datasets. This is expected since the $\ell_{\infty}$ values for bounded attacks are constrained by the attack budget settings outlined in Section 4.1. For C&W and DeepFool, the lowest $\ell_{\infty}$ values are achieved in most cases, except for the German dataset. Similar to the observation on proximity $\ell_{2}$, the C&W attack yields the lowest $\ell_{\infty}$ values on both LR and MLP models, while DeepFool consistently outperforms on the LinearSVC model. In contrast, LowProFool produces the highest $\ell_{\infty}$ values among unbounded attacks on the Diabetes and Breast Cancer datasets.
@@
Deviation metric assesses how perturbed examples relate to the overall dataset distribution. When the deviation metric is notably high, adversarial attacks are more likely to produce examples outside of the dataset’s distribution. Therefore, lower deviation values are preferable for crafting imperceptible attacks. In Table 8, the evaluation of the deviation metric reveals that C&W consistently yields the lowest deviation value for both LR and MLP models, implying a higher degree of similarity to the corresponding input dataset’s statistical distribution. When attacking the LinearSVC model, DeepFool produces the lowest deviation value. The deviation value of LowProFool is higher than that of C&W and DeepFool, but lower than the two bounded attacks. In contrast, the two bounded attacks, FGSM and PGD, share the similar deviation values and consistently yield higher deviation values, suggesting that they tend to produce more out-of-distribution data.
Table 8: Results of deviation of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models over five datasets. The best deviation for each model is shown in bold. Due to insufficient effectiveness, the combination of C&W attack and LinearSVC is not considered and their results are marked with $\dagger$.
@@
Sensitivity quantifies how much an attack algorithm perturbs those features with low variability (i.e. features with a narrow distribution). Lower sensitivity values are preferable because attack algorithms would introduce fewer perturbation on features with a narrow distribution, which in turn reduces imperceptibility. As shown in Table 9, for each individual dataset, two bounded attacks generally exhibit similar values in deviation, which are higher than those observed for C&W and DeepFool attacks. However, LowProFool, an unbounded attack, shows similar results to FGSM and PGD. Regarding LR and MLP models, the C&W attack demonstrates lowest values in sensitivity metric with exception of the combination of German dataset and LR model. DeepFool always yeilds the lowest sensitivity values in all five dataset on LinearSVC. This indicates that LowProFool, FGSM, and PGD attacks tend to introduce more significant perturbations on narrow distribution features than C&W and DeepFool.
Table 9: Results of sensitivity of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models over five datasets. The best sensitivity for each model is shown in bold. Due to insufficient effectiveness, the combination of C&W attack and LinearSVC is not considered and their results are marked with $\dagger$.
@@
Summary
The evaluation of the four quantitative imperceptibility properties suggests that overall the three unbounded attacks, C&W, DeepFool and LowProFool, generate more imperceptible adversarial examples compared to the two bounded attacks, FGSM and PGD. C&W often achieves the best imperceptibility on LR and MLP across different datasets, and DeepFool yields the best imperceptibility on LinearSVC. In contrast, LowProFool demonstrates the worst imperceptibility among three unbounded attacks in our evaluation. The two bounded attacks, FGSM and PGD, exhibit similar results across each property evaluation, likely due to their shared attack mechanism, differing only in their one-time versus iterative application.
@@ 4.4 Analysis of Imperceptibility using Qualitative Properties
In Section 3, we define three qualitative properties of attack imperceptibility, which are Immutability, feasibility and feature interdependency, and they often rely on domain-specific knowledge about the datasets. The systematic process for analysing and evaluating adversarial examples based on these qualitative properties involves the following key points:
Step 1: Analyse and Identify Key Features
  • 1.
    Immutability: Identify features that must remain unchanged during any modification. This requires a thorough understanding of domain-specific knowledge. For example, certain attributes like gender and age must remain unchanged in the context of the dataset.
  • 2.
    Feasibility: Define the feasible range for each modifiable feature, ensuring the values stay realistic within the data domain. This can include upper and lower bounds, categorical constraints, or logical limitations specific to the dataset. While statistical methods can be employed to identify frequent and infrequent values, determining whether infrequent values are truly infeasible requires relevant domain knowledge.
  • 3.
    Feature Interdependency: Analyse the relationships or correlations between different features, ensuring modifications respect these dependencies. For instance, an increase in one feature may necessitate a corresponding increase or decrease in another, depending on the domain-specific rules. While statistical methods can reveal correlations between individual features, domain knowledge is essential to identify and understand more complex interdependencies.
Step 2: Evaluate Adversarial Examples Against Properties
  • 1.
    Immutability: Does the attack alter features that should remain unchanged?
  • 2.
    Feasibility: Are the modifications within the feasible ranges of the corresponding features?
  • 3.
    Feature Interdependency: Are the feature interdependencies between features preserved after the attack?
Our current evaluation prioritises the use of domain knowledge to assess the qualitative properties of immutability, feasibility, and feature interdependency. Statistical techniques, such as identifying common value ranges and detecting correlations between features, can offer an additional layer of objectivity and precision. These methods will be explored and incorporated in future work.
In our current work, for immutability, we applied common knowledge as domain knowledge and have checked all the adversarial examples generated by each of the evaluated attacks on all three datasets containing mixed variables (i.e. Adult, COMPAS, and German). For feasibility, we used the collected domain knowledge (see Table 15) to identify the feasible value ranges for each feature in the Diabetes dataset and evaluated all the adversarial attacks to check if any perturbed features fall outside these ranges. We also observed apparent feature interdependencies in the COMPAS dataset and evaluated whether the generated adversarial examples violate these interdependencies.
When presenting the evaluation results, it is impractical to explain the results for all three properties across every combination of dataset, model, and attack. Therefore, we have selected a few representative examples to illustrate the key findings on immutability, feasibility, and feature interdependency.
@@
We analyse immutability across three mixed datasets—Adult, COMPAS and German, which we can identify examples of immutable features using common knowledge. These examples of immutable features in each of these datasets are shown below:
@@
    German: Personal Status & Sex and Foreign Worker Status
Figure 3: Heatmap showing the percentage of adversarial examples generated for each combination of model (LR, LinearSVC, MLP) and attack type (DeepFool, C&W, FGSM, PGD) that perturb immutable features: race, sex, and marital-status. The values are presented as a percentage of a total of 6513 adversarial examples.
We examine each adversarial example to check whether the above immutable feature values have been modified. This allows us to assess whether the identified immutable features are susceptible to manipulation. The results of perturbing immutable features for each of the three datasets are detailed in A.1. We observe that DeepFool successfully perturbs immutable features in both LR and MLP models across all three datasets, violating the immutability requirement. For instance, in the Adult dataset (Figure 3), DeepFool generates a total of 6,513 adversarial examples when attacking the MLP model. In these examples, the feature Race is modified in 648 instances (9.95%), Sex in 345 instances (5.3%), and Marital Status in 1,096 instances (16.83%).
Conversely, C&W, FGSM and PGD do not appear to alter these immutable features. However, this does not necessarily mean that the three attacks comply with the immutability requirement. Theoretically, the design of these attacks does not consider any strategy for avoiding perturbation of immutable features. A close examination of these attacks has revealed that they are incapable of changing immutable features—which are all categorical features in the examples used and are preprocessed with one-hot encoding in implementation.
To investigate how immutable features are perturbed by DeepFool, we use the COMPAS dataset and LR model as an example to illustrate how DeepFool perturbs immutable features. The COMPAS dataset, known for its racial bias in predicting the likelihood of recidivism either as “Medium-Low” or “High” risk (Barenstein, 2019). In our experiment, race and sex are considered as immutable features. We generate 1,408 adversarial examples for DeepFool on the LR model, and retrieve 8 cases with perturbation on race (0 case for sex). Table 10 provides two example cases (#285 and #501) out of these 8 cases, which are correctly predicted as “Medium-Low” by the LR model. DeepFool successfully misclassified the predictions of these two cases by altering the value of the immutable feature race from “Caucasian” to “Native American” (Case #285) and from “Hispanic” to “African-American” (Case #501), respectively.
Table 10: Partial feature values between original datapoints and adversarial examples generated by DeepFool attack on LR in Case #285 and Case #501 over the COMPAS dataset.
@@
#285 Original 80 0 0 Greater than 45 Male Caucasian Medium-Low
@@
#501 Original 83 0 0 Greater than 45 Male Hispanic Medium-Low
@@
We also examine the feature weights (coefficients) of the LR model, as depicted in Figure 4. Considering the two example cases in Table 10, DeepFool tends to perturb features with considerable impacts on model prediction, and race is among these features. In contrast, the other immutable feature, sex, has little impact on model prediction, and therefore it is not modified by DeepFool.
Figure 4: Feature weights (coefficients) of LR for the COMPAS dataset. The features are ranked by feature weights. Features with a positive weight contribute to “Medium-Low” risk and features with a negative weight contribute to “High” risk. Features with a weight in the range of $[-0.15,0.15]$ have little impact on model prediction and therefore are omitted.
@@
This property assesses if adversarial attack algorithms can perturb feature values within a reasonable range. However, since there is a shortage of background knowledge for all datasets, it is impractical to identify all feasible feature ranges for each dataset. To assess the feasibility of adversarial attacks, we turn to the Diabetes dataset. It only comprises numerical features capturing diagnostic measurements which are used to predict diabetes. According to the description of each variable in the Diabetes dataset (Smith et al., 1988), we identify feasible value ranges for all features and these are listed in A.2.
Table 11: Partial feature values between original datapoints and adversarial examples generated by DeepFool, C&W and LowProFool attacks in Case #19 and Case #57 over the Diabetes dataset.
@@
To investigate how adversarial attacks modify features beyond their feasible ranges, we select two case examples #19 and #57 from the generated adversarial examples with the LR model. We then compare the original feature values with the adversarial examples generated by DeepFool, C&W, and LowProFool attack in Table 11. The following observations can be made:
@@
    Glucose: In both cases, all three attack methods increase the value of Glucose from below 90 to above 135. The C&W attack introduces comparatively minor perturbations, and the revised Glucose values are still within the normal range (below 140). The DeepFool and LowProFool attacks introduce more substantial changes, pushing the glucose levels into the impaired glucose tolerance range (140–199). Particularly with LowProFool, it alters the Glucose value to reach 199 in both cases, which is extremely close to the glucose threshold (200) indicating diabetes.
@@
    SkinThickness: The adversarial examples generated by DeepFool in Case #57 and by LowProFool in both cases exceed the maximum reference value (23.6$\pm$7.5mm).
@@
    Insulin: In Case #57, both the DeepFool and LowProFool attacks decrease the insulin level to below the lower bound of the recommended 2-hour serum insulin range (16–166). This perturbed feature value is not feasible regarding the expected physiological behavior.
@@
    BMI: In both cases, all attacks modify the BMI values from around 30 to over 40, indicating severe obesity. Particularly in Case #57, the DeepFool and LowProFool attacks perturb the BMI to 59.85 and 67.10, respectively, which are unrealistic.
Figure 5: Feature weights (coefficients) of LR for the Diabetes dataset. The features are ranked by feature weights. Features with a positive weight contribute to “has” diabetes and features with a negative weight contribute to “not have” diabetes.
The feature weights of the LR model for the Diabetes dataset are visualised in Figure 5. It is evident that Glucose and BMI are the top two most important features contributing to the prediction of having diabetes. This explains why all the above three attacks tend to introduce significant perturbations in these two features, often pushing their resulting feature values into infeasible ranges.